What is the percentage concentration of 150.0g of ethanol dissolved in 500.0g of water
1
Expert's answer
2020-04-09T13:11:06-0400

С= m(C2H5OH)/m(solution) * 100%=150/650 * 100%=23,08%

Answer: 23.08%